Hi
Joe and Bob took part in a race. When Bob had completed the race in 30min, Joe ran only 4/7 of the distance. Joe's average speed for the race was 80m per minute less than Bob's average speed.
Find the distance of the race.
What was Joe's speed in km per hour.
Thanks

Found 2 solutions by 1931, MathTherapy:
Answer by 1931(1)

Let the avg speed of Joe is x km/hr then avg speed of Bob = x+80 km/hr
and the d is the distance of race
Bob had completed the race in 30 min
Thus d=(x+80)/2..............(1)
Distance covered by Joe in 30 min
= x/2 =4/7*d
x/2= 4/7* (x+80)/2
x/2=2(x+80)/7
7x=4(x+80)
7x=4x+320
3x=320
x=320/3
Thus Joe's speed =106.66 km/hr
From(1)
d=(106.66+80)/2
=186.66/2=93.33 km
Thus Distance of race= 93.33 Km



Answer by MathTherapy(10810)


Hi
Joe and Bob took part in a race. When Bob had completed the race in 30min, Joe ran only 4/7 of the distance. Joe's average speed for the race was 80m per minute less than Bob's average speed.
Find the distance of the race.
What was Joe's speed in km per hour.
Thanks
That person is WAY, WAY off! Which human has been known to run over 100 km/h? RIDICULOUS!!